# The geometry of preperiodic points in families of maps on $\mathbb{P}^{N}$

### v1

@@ 3.4. Common preperiodic points in $\mathbb{P}^{1}$ and uniform bounds
It was proved in [BD1] and [YZ2] that any pair of maps $f,g:\mathbb{P}^{1}\to\mathbb{P}^{1}$ over $\mathbb{C}$, of degrees $>1$, will satisfy either $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ or $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|<\infty$. This result can be viewed as a special case of Conjecture 1.1, at least when $\deg f=\deg g$. Indeed, suppose that $\deg(f)=\deg(g)=d\geq 2$, and consider the action of $\Phi:=(f,g)$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$. This $\Phi$ is polarizable and so extends to an endomorphism of some $\mathbb{P}^{N}$, restricting to the given map $(f,g)$ on an embedded copy of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The common preperiodic points in $\mathbb{P}^{1}$ for $f$ and $g$ correspond to preperiodic points of $\Phi$ in the diagonal $\Delta\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$. Conjecture 1.1 reduces to the DMM conjecture in this setting, as discussed in §3.1; explicitly, it states that $\Delta$ is $\Phi$-special if and only if $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|=\infty$. The equivalence of equality $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ with $\Delta$ being $\Phi$-special follows from [YZ2, Theorem 1.4] combined with [LP, Theorem A] and [MS, Theorem 1.10].
Conjecture 1.1 also predicts the existence of uniform bounds on the size of $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|$ in families of pairs. For example, it was proved in [DM] that there exists a Zariski-open and -dense subset $U_{d}$ of the space of all pairs $\mathrm{Rat}_{d}\times\mathrm{Rat}_{d}$ of degree $d$, and a constant $B_{d}$, so that
(3.2) $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|\leq B_{d}$
for all $(f,g)\in U_{d}$; this theorem can be seen as a special case of Conjecture 1.1 (and in fact of Conjecture 1.2), as follows.
@@
By [DM, Lemma 5.4], to prove (3.2) it suffices to show that there exists $k\geq 1$ so that $\mathcal{X}_{k}$ does not contain a Zariski dense set of $\Phi^{[k]}$-preperiodic points. We claim that the relative special dimension $r_{\Phi^{[k]},\mathcal{X}_{k}}$ is equal to $2k$ for any $k\geq 1$; that is, the subvariety $\mathcal{X}_{k}$ of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ is not contained in any $\Phi^{[k]}$-special subvariety other than $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ itself. Then (3.2) follows immediately from Conjecture 1.2, for any $k>\dim S_{d}$.
Let $\mathcal{Z}\subset S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ be ${\Phi}^{[k]}$-special and assume the generic fiber $\mathbf{Z}$ contains $\mathbf{X}_{k}$. Since there exists a pair of non-exceptional maps $(f_{s},g_{s})$ for $s\in S_{d}$ with distinct Julia sets, it must be that $\mathbf{Z}$ is $\mathbf{\Phi}^{[k]}$-preperiodic and neither $\mathbf{f}$ nor $\mathbf{g}$ is exceptional; this follows from the structural result on invariant subvarieties for product maps on $(\mathbb{P}^{1})^{N}$ (for any $N$) over fields of characteristic 0, proved in [MeSc] (and reproved by different methods in [GNY]). It follows from this classification that $\mathcal{Z}$ must be all of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$. This completes the proof.
@@
Conjecture 4.1 fails if we do not assume that $S\to\mathrm{M}_{d}^{N}$ is dominant. For example, $\Phi$ could be a family of regular polynomial maps on $\mathbb{P}^{2}$ (with coordinates $(x:y:z)$), where the line at infinity $\{z=0\}$ is invariant for $\Phi_{s}$, for all $s\in S$. Even if we further assume that $\mathcal{X}$ is not $\Phi$-special, the conjecture can fail without assuming dominance of $S\to\mathrm{M}_{d}^{N}$; for example, the family $\mathcal{X}$ of lines $\{x=s\,y\}$ in $\mathbb{P}^{2}$, for $s\in\mathbb{C}$, is not special under $\Phi(s,x,y)=(s,x^{2},y^{2})$ but becomes $\Phi_{s}$-preperiodic for each root of unity $s$.
@@
As for the example of §3.4, we consider fiber powers of the family $\Phi$. Let
@@
for some large $D(N,m)$. The subvariety $S\times(\mathbb{P}^{N})^{m}$ will sit inside $S\times\mathbb{P}^{D(N,m)}$ as a $\Phi_{N,m}$-invariant subvariety, so we will restrict our attention to $\Phi^{[m]}$ itself.
Let $\mathcal{X}_{m}$ denote the $m$-th fiber power of $\mathcal{X}$ over $S$, as a subvariety of $S\times(\mathbb{P}^{N})^{m}$, so each fiber of the projection $\mathcal{X}_{m}\to S$ is of the form $X_{s}^{m}$. Let $r_{m}^{N}$ be the relative special dimension of $\mathcal{X}_{m}$, so that
$r_{m}^{N}:=r_{\Phi^{[m]},\mathcal{X}_{m}}\leq mN=\dim\,(\mathbb{P}^{N})^{m}.$
@@
$\dim\mathcal{X}_{m}\geq r_{m}^{N}$
for all $m\geq 1$. We will see, however, that $r_{m}^{N}=mN$ for all $m$, so that if $m>\dim S$, then
$r_{m}^{N}>\dim\mathcal{X}_{m}=\dim S+m\dim_{S}(\mathcal{X}),$
@@
To see that $r_{m}^{N}=mN$, we work inductively on $m$ and appeal to a result of Fakhruddin [Fa2, Theorem 1.2]. He proved that any Zariski-closed subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ which projects dominantly to $S$ and is invariant for $\Phi$ must either be all of $S\times\mathbb{P}^{N}$ or have codimension $N$. Note that there are no polarizable endomorphisms $\Psi$ commuting with $\Phi$ over all of $S$, except the iterates of $\Phi$ itself, so $\Phi$-special is the same as $\Phi$-preperiodic in this setting. Moreover, any Zariski-closed $\Phi$-invariant $\mathcal{Y}$ which projects dominantly to $S$ and has codimension $N$ must parameterize a finite collection of preperiodic points for $\Phi$; in particular, such a $\mathcal{Y}$ must project surjectively to $S$ and the fiber over every parameter in $S$ will have codimension $N$ in $\mathbb{P}^{N}$.
For $m=1$, we have $\mathcal{X}_{1}:=\mathcal{X}$ and by [Fa2, Theorem 1.2] the only $\Phi$-special subvariety containing $\mathcal{X}$ is $S\times\mathbb{P}^{N}$. So $r_{1}^{N}=N$. Now fix $m>1$, and assume that $r_{j}^{N}=jN$ for all $j<m$. Suppose that $\mathcal{Z}_{m}$ is a $\Phi^{[m]}$-special subvariety in $S\times(\mathbb{P}^{N})^{m}$ containing $\mathcal{X}_{m}$. We aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$.
@@
denote the projection forgetting the last coordinate $\mathbb{P}^{N}$. By the induction hypothesis, we have $p(\mathcal{Z}_{m})=S\times(\mathbb{P}^{N})^{m-1}$, because $p(\mathcal{Z}_{m})$ must be $\Phi^{[m-1]}$-special and contain $\mathcal{X}_{m-1}$. Let $\mathcal{P}_{1}\subset S\times(\mathbb{P}^{N})^{m-1}$ be a subvariety of dimension $=\dim S$ which projects surjectively to $S$ and is pointwise fixed by $\Phi^{[m-1]}$. (In other words, $\mathcal{P}_{1}$ is a multisection of the projection from $S\times(\mathbb{P}^{N})^{m-1}$ to $S$, parameterizing a collection of fixed points.) Consider
@@
where $x$ is an element of $\mathcal{P}_{1}$, and $p_{m}:S\times(\mathbb{P}^{N})^{m}\to\mathbb{P}^{N}$ is the projection to the last coordinate. Note that
@@
for every $x\in\mathcal{P}_{1}$ and $y\in\mathcal{Z}^{x}_{m}$, because $x$ is a fixed point. Since $\mathcal{Z}_{m}$ is preperiodic under $\Phi^{[m]}$, we see that $\mathcal{Z}^{x}_{m}$ defines a family of $\Phi$-preperiodic subvarieties in $\mathbb{P}^{N}$, as $x$ varies over $\mathcal{P}_{1}$. It follows from Fakhruddin’s theorem [Fa2, Theorem 1.2] that $\{\mathcal{Z}^{x}_{m}:x\in\mathcal{P}_{1}\}$ is either a family of points or all of $\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$.
@@
In the former case, we deduce that $\operatorname{codim}\mathcal{Z}_{m}=N$ in $S\times(\mathbb{P}^{N})^{m}$, because the dimension of the fibers of $p|_{\mathcal{Z}_{m}}$ is upper semi-continuous over $S\times(\mathbb{P}^{N})^{m-1}$. It follows that $\mathcal{Z}_{m}$ defines a family of preperiodic points in $\mathbb{P}^{N}$ over all of $S\times(\mathbb{P}^{N})^{m-1}$ (viewed as an extended parameter space for the family of maps $\Phi_{s}$). But the fibers of the projection $p|_{\mathcal{Z}_{m}}$ over $\mathcal{X}_{m-1}$ must contain the corresponding fibers of $\mathcal{X}$, a contradiction.
@@
We conclude that $\mathcal{Z}^{x}_{m}=\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$. Recall that we aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$. We now repeat this argument, replacing $\mathcal{P}_{1}$ with any family $\mathcal{P}$ of marked periodic points for $\Phi^{[m-1]}$ in $S\times(\mathbb{P}^{N})^{m-1}$ and work with an iterate of $\Phi$, and we deduce that the fibers of $\mathcal{Z}_{m}$ over $\mathcal{P}$ are also all of $\mathbb{P}^{N}$. As the periodic points of $\Phi^{[m-1]}$ are Zariski-dense in $S\times(\mathbb{P}^{N})^{m-1}$, and as the fibers of $p|_{\mathcal{Z}_{m}}$ are equal to $\mathbb{P}^{N}$ over this Zariski-dense set, we conclude that every fiber of $p|_{\mathcal{Z}_{m}}$ is equal to $\mathbb{P}^{N}$. Therefore, $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$ and so the relative special dimension of $\mathcal{X}_{m}$ is $r_{m}^{N}=mN$. ∎
4.1. PCF density and sparsity in the moduli space

### v2

@@ 3.4. Common preperiodic points in $\mathbb{P}^{1}$ and uniform bounds
For a map $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ let $\mathrm{Prep}(f)\subset\mathbb{P}^{1}(\mathbb{C})$ denote its set of preperiodic points. It was proved in [BD1] and [YZ2] that any pair of maps $f,g:\mathbb{P}^{1}\to\mathbb{P}^{1}$ over $\mathbb{C}$, of degrees >1>1, will satisfy either $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ or $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|<\infty$. This result can be viewed as a special case of Conjecture 1.1, at least when $\deg f=\deg g$. Indeed, suppose that $\deg(f)=\deg(g)=d\geq 2$, and consider the action of $\Phi:=(f,g)$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$. This $\Phi$ is polarizable and so extends to an endomorphism of some $\mathbb{P}^{N}$, restricting to the given map $(f,g)$ on an embedded copy of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The common preperiodic points in $\mathbb{P}^{1}$ for $f$ and $g$ correspond to preperiodic points of $\Phi$ in the diagonal $\Delta\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$. Conjecture 1.1 reduces to the DMM conjecture in this setting, as discussed in §3.1; explicitly, it implies that $\Delta$ is $\Phi$-special if and only if $|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|=\infty$. The equivalence of the equality $\mathrm{Prep}(f)=\mathrm{Prep}(g)$ with $\Delta$ being $\Phi$-special follows from [YZ2, Theorem 1.4] combined with [LP, Theorem A] and [MS, Theorem 1.10].
Conjecture 1.1 (or its weaker form, Conjecture 1.3) predicts what we view as uniform versions of itself. For example, in this setting of shared preperiodic points for two maps on $\mathbb{P}^{1}$, it implies the existence of the uniform bound that we proved in [DM, Theorem 1.1]:
Proposition 3.6.
Conjecture 1.3 implies that for each degree $d\geq 2$, there exists a constant $B_{d}>0$ and a Zariski-open subset $U_{d}$ of the space $\mathrm{Rat}_{d}\times\mathrm{Rat}_{d}$ of all pairs of maps $f,g:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree $d$ so that
$|\mathrm{Prep}(f)\cap\mathrm{Prep}(g)|\leq B_{d}$
for all $(f,g)\in U_{d}$.
Proof.
@@
In [DM, Lemma 5.4], we observed that if the conclusion of the proposition were to fail, then $\mathcal{X}_{k}$ will contain a Zariski dense set of $\Phi^{[k]}$-preperiodic points for all $k\geq 1$. So it suffices to show that there exists $k\geq 1$ so that $\mathcal{X}_{k}$ does not contain a Zariski dense set of $\Phi^{[k]}$-preperiodic points.
To this end, we will show that the relative special dimension $r_{\Phi^{[k]},\mathcal{X}_{k}}$ is equal to $2k$ for any $k\geq 1$; that is, the subvariety $\mathcal{X}_{k}$ of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ is not contained in any $\Phi^{[k]}$-special subvariety other than $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ itself. Then, taking any $k>\dim S_{d}$, we obtain our desired conclusion from Conjecture 1.3.
Let $\mathcal{Z}_{k}\subset S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$ be an irreducible ${\Phi}^{[k]}$-special variety that contains $\mathcal{X}_{k}$. It must be that $\mathcal{Z}_{k}$ is preperiodic for $\Phi^{[k]}$, as the map $\Phi_{s}$ commutes with nothing but its iterates for a generic choice of $s\in S_{d}$; see for example [Ye, Theorem 1.2]. From the structural results on invariant subvarieties for product maps on $(\mathbb{P}^{1})^{N}$ (for any $N$) over fields of characteristic 0, proved in [MeSc] (and reproved by different methods in [GNY2]), it follows that $\mathcal{Z}$ must be all of $S_{d}\times(\mathbb{P}^{1}\times\mathbb{P}^{1})^{k}$. That is, $r_{\Phi^{[k]},\mathcal{X}_{k}}=2k$, and the proof is complete. ∎
@@
Conjecture 4.1 fails if we do not assume that $S\to\mathrm{M}_{d}^{N}$ is dominant. For example, $\Phi$ could be a family of regular polynomial maps on $\mathbb{P}^{2}$ (with coordinates $(x:y:z)$), where the line at infinity $\mathcal{X}=\{z=0\}$ is invariant for $\Phi_{s}$, for all $s\in S$. Even if we further assume that $\mathcal{X}$ is not $\Phi$-special, the conjecture can fail without assuming dominance of $S\to\mathrm{M}_{d}^{N}$; for example, the family $\mathcal{X}$ of lines $\{x=s\,y\}$ in $\mathbb{P}^{2}$, for $s\in\mathbb{C}$, is not special under $\Phi(s,(x:y:z))=(s,(x^{2}:y^{2}:z^{2}))$ but becomes $\Phi_{s}$-preperiodic for each root of unity $s$.
@@
As in the proof of Proposition 3.6, we consider fiber powers of the family $\Phi$. Let
@@
for some large $D(N,m)$, restricting to a Zariski-open subset of $S$ if necessary. The subvariety $S\times(\mathbb{P}^{N})^{m}$ will sit inside $S\times\mathbb{P}^{D(N,m)}$ as a $\Phi_{N,m}$-invariant subvariety, so we will restrict our attention to $\Phi^{[m]}$ itself.
Let $\mathcal{X}_{m}$ denote the $m$-th fiber power of $\mathcal{X}$ over $S$, as a subvariety of $S\times(\mathbb{P}^{N})^{m}$, so each fiber of the projection $\mathcal{X}_{m}\to S$ is of the form $X_{s}^{m}$. Let $r_{N,m}$ be the relative special dimension of $\mathcal{X}_{m}$, so that
$r_{N,m}:=r_{\Phi^{[m]},\mathcal{X}_{m}}\leq mN=\dim\,(\mathbb{P}^{N})^{m}.$
@@
$\dim\mathcal{X}_{m}\geq r_{N,m}$
for all $m\geq 1$. We will see, however, that $r_{N,m}=mN$ for all $m$, so that if $m>\dim S$, then
$r_{N,m}>\dim\mathcal{X}_{m}=\dim S+m\dim_{S}(\mathcal{X}),$
@@
To see that $r_{N,m}=mN$, we work inductively on $m$ and appeal to a result of Fakhruddin [Fa2, Theorem 1.2]. He proved that any irreducible Zariski-closed subvariety $\mathcal{Y}\subset S\times\mathbb{P}^{N}$ which projects dominantly to $S$ and is invariant for $\Phi$ must either be all of $S\times\mathbb{P}^{N}$ or have codimension $N$. Moreover, if $\mathcal{Y}$ has codimension $N$, then it must parameterize a finite collection of preperiodic points for $\Phi$. Note that there are no polarizable endomorphisms $\Psi$ commuting with $\Phi$ over all of $S$, except the iterates of $\Phi$ itself, so $\Phi$-special is the same as $\Phi$-preperiodic in this setting.
For $m=1$, we have $\mathcal{X}_{1}:=\mathcal{X}$ and by [Fa2, Theorem 1.2] the only $\Phi$-special subvariety containing $\mathcal{X}$ is $S\times\mathbb{P}^{N}$. So $r_{N,1}=N$. Now fix $m>1$, and assume that $r_{N,j}=jN$ for all $j<m$. Suppose that $\mathcal{Z}_{m}$ is a $\Phi^{[m]}$-special subvariety in $S\times(\mathbb{P}^{N})^{m}$ containing $\mathcal{X}_{m}$. We aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$.
@@
denote the projection forgetting the last coordinate $\mathbb{P}^{N}$. By the induction hypothesis, we have $p(\mathcal{Z}_{m})=S\times(\mathbb{P}^{N})^{m-1}$, because $p(\mathcal{Z}_{m})$ must be $\Phi^{[m-1]}$-special and contain $\mathcal{X}_{m-1}$. Let $\mathcal{P}_{1}\subset S\times(\mathbb{P}^{N})^{m-1}$ be a subvariety of dimension $\dim S$ which projects surjectively to $S$ and is pointwise fixed by $\Phi^{[m-1]}$. (In other words, $\mathcal{P}_{1}$ is a multisection of the projection from $S\times(\mathbb{P}^{N})^{m-1}$ to $S$, parameterizing a collection of fixed points.) For each $x\in\mathcal{P}_{1}$, consider
@@
where $p_{m}:S\times(\mathbb{P}^{N})^{m}\to\mathbb{P}^{N}$ is the projection to the last coordinate. Note that
@@
for every $x\in\mathcal{P}_{1}$ and $y\in\mathcal{Z}^{x}_{m}$, because $x$ is a fixed point. As $\mathcal{Z}_{m}$ is preperiodic under $\Phi^{[m]}$, and letting $\Phi_{\mathcal{P}_{1}}$ denote the family $\Phi$ of maps on $\mathbb{P}^{N}$ over $\mathcal{P}_{1}$ via the base change $\mathcal{P}_{1}\to S$, we see that $\{\mathcal{Z}^{x}_{m}:x\in\mathcal{P}_{1}\}$ defines a family of $\Phi_{\mathcal{P}_{1}}$-preperiodic subvarieties in $\mathbb{P}^{N}$. Applying Fakhruddin’s theorem [Fa2, Theorem 1.2] to each irreducible component of this family, we see that $\{\mathcal{Z}^{x}_{m}:x\in\mathcal{P}_{1}\}$ is either a family of points or all of $\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$.
@@
In the former case, we deduce that $\operatorname{codim}\mathcal{Z}_{m}=N$ in $S\times(\mathbb{P}^{N})^{m}$, because the dimension of the fibers of $p|_{\mathcal{Z}_{m}}$ is upper semi-continuous over $S\times(\mathbb{P}^{N})^{m-1}$. Again applying [Fa2, Theorem 1.2], the irreducibility of $\mathcal{Z}_{m}$ means that it defines a family of preperiodic points in $\mathbb{P}^{N}$ over all of $S\times(\mathbb{P}^{N})^{m-1}$ (viewed as an extended parameter space for the family of maps $\Phi_{s}$ via the projection $p$). But the fibers of the projection $p|_{\mathcal{Z}_{m}}$ over $\mathcal{X}_{m-1}$ must contain the corresponding fibers of $\mathcal{X}$, a contradiction.
@@
We conclude that $\mathcal{Z}^{x}_{m}=\mathbb{P}^{N}$ for all $x\in\mathcal{P}_{1}$. Recall that we aim to show that $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$. We now repeat this argument, replacing $\mathcal{P}_{1}$ with any family $\mathcal{P}$ of marked periodic points for $\Phi^{[m-1]}$ in $S\times(\mathbb{P}^{N})^{m-1}$ and work with an iterate of $\Phi$, and we deduce that the fibers of $\mathcal{Z}_{m}$ over $\mathcal{P}$ are also all of $\mathbb{P}^{N}$. As the periodic points of $\Phi^{[m-1]}$ are Zariski-dense in $S\times(\mathbb{P}^{N})^{m-1}$, and as the fibers of $p|_{\mathcal{Z}_{m}}$ are equal to $\mathbb{P}^{N}$ over this Zariski-dense set, we conclude that every fiber of $p|_{\mathcal{Z}_{m}}$ is equal to $\mathbb{P}^{N}$. Therefore, $\mathcal{Z}_{m}=S\times(\mathbb{P}^{N})^{m}$ and so the relative special dimension of $\mathcal{X}_{m}$ is $r_{N,m}=mN$. ∎
4.1. PCF density or sparsity in the moduli space